# Uncertainty Quantification on Graph Learning: A Survey

### v2

@@ V-A1 Distributionally Robust Optimization
Distributionally Robust Optimization (DRO) aims to enhance the model’s performance across a wide range of potential testing data distributions, instead of the training data distribution only. Specifically, DRO optimizes the objective function in the “worst case” of the distributions where the loss will be maximized. Formally, DRO is formulated by a bi-level optimization problem:
$\min_{f\in\mathcal{F}}\max_{P\in\mathcal{P}}\mathbb{E}_{x\sim P}[\mathcal{L}_{0}(f,x)],$ (26)
where $\mathcal{L}_{0}$ is a classification loss, such as cross-entropy. The uncertainty set $\mathcal{P}:=\{P|D(P,P_{0})\leq\rho\}$ includes all the data distributions $P$ within the radius distance $\rho$ from the observed data distribution $P_{0}$. The distance between two distributions is measured by $D(\cdot,\cdot)$, and we list the following choices of $D$ used in DRO.
  • $\mathbf{f}$-divergence, also known as phi-divergence, is defined by $D_{f}(P||P_{0}):=\int f\left(\frac{dP}{dP_{0}}\right)dP_{0}$. It is widely used to quantify the divergence between the distributions in DRO [90, 91, 92]. The domain delineated by $f$-divergence can be viewed as a statistical confidence region, and the inner maximization problem in Eq. (26) is tractable for several choices of $f$ [92, 93].
  • KL-divergence, one special case of $f$-divergence, is used as a distance metric in DRO [94, 95]. However, the possible limitations of metrics based on $f$-divergence are that they may not be comprehensive enough to include some relevant distributions, or they may fail to precisely measure the distance for extreme distributions [96].
[…]
  • Besides that, authors in [102] study DRO with uncertainty sets defined by maximum mean discrepancy. The authors in [103] propose to define the uncertainty sets using the Prohorov metric.
V-A2 Adversasrial Training and Self-Supervised Learning
Besides DRO, adversarial training (AT) and self-supervised learning (SSL) are proposed to mitigate OOD problems. However, most of the existing work adopts the $L_{p}$-norm or predefined perturbations to quantify or mimic the uncertainty.
In essence, AT on graphs addresses the following min-max optimization problem:
$\min_{\boldsymbol{\theta}}\max_{\|\epsilon_{\theta}\|\leq\rho_{\theta},\|\epsilon_{x}\|\leq\rho_{x},\|\epsilon_{a}\|\leq\rho_{a}}\mathcal{L}_{0}(\boldsymbol{\theta}+\epsilon_{\theta},X+\epsilon_{x},A+\epsilon_{a}),$ (27)
where $\epsilon_{\theta}$, $\epsilon_{x}$, and $\epsilon_{a}$ represent the perturbations applied to model parameters $\boldsymbol{\theta}$, input node features $X$, and adjacency matrix $A$, respectively. $\rho_{\theta}$, $\rho_{x}$ and $\rho_{a}$ ensures the perturbations $\epsilon_{\theta}$, $\epsilon_{x}$, and $\epsilon_{a}$ to be small enough and neglectable, respectively. Notice that the perturbation magnitude (uncertainty) is typically measured using the $L_{p}$-norm, such as the $L_{1}$, $L_{2}$, or $L_{\infty}$ norm, regardless of whether the input data [104, 106] or the model parameters [106, 105] are being perturbed.
The idea of SSL on graphs is to deliberate perturbation or modification of graphs to construct a new controllable dataset in which the intrinsic information of the graphs remains unchanged. The modifications are typically predefined and explicitly formulated, thus these works cannot provide precise measurements for changes (uncertainty). To facilitate the generation of positive pair instances, these works encompass augmentations such as node masking [107], edge perturbation, attribute masking, subgraph sampling [108], random walks for graph sampling [109], and both structural and attribute levels augmentations [110].
V-B Conformal Prediction
V-B1 Conformal Prediction Overview
Conformal Prediction (CP) is an uncertainty quantification framework. For a given model and a predefined significant level $\alpha$, CP uses the calibration data to estimate the confidence intervals of predicted outcomes based on score functions [111]. In this section, we focus on CP for classification tasks, as GNN-related tasks inherently fall into this category.

### v3

@@ V-B1 Distributionally Robust Optimization
Distributionally Robust Optimization Overview. Distributionally Robust Optimization (DRO) aims to enhance the model’s performance across a wide range of potential testing data distributions, instead of the training data distribution only. Specifically, DRO optimizes the objective function in the “worst case” of sample distributions $P\in\mathcal{P}$, in which place will the loss be maximized. Eq. (32) highlights the objective of DRO, which is formulated as a bi-level optimization problem:
$\min_{\boldsymbol{\theta}}\max_{P\in\mathcal{P}}\mathbb{E}_{(x,y)\sim P}[\mathcal{L}_{0}(\boldsymbol{\theta},x,y)],$ (32)
where $\mathcal{L}_{0}$ is a loss function. The uncertainty set $\mathcal{P}$, formally defined as $\mathcal{P}:=\{P|D(P,P_{0})\leq\rho\}$, encapsulates all the unseen data distribution within a specific distance from the training distribution. As discussed in Sec. IV-C1, several divergence metrics $D$ have been adopted in DRO.
Distributionally Robust Optimization for Graphs. Due to the unique topological structure of the graphs and various tasks involved, such as node classification and graph classification, the distributions could be defined on node and graph level, respectively. Specifically, the distance metric $D$ is employed to evaluate the discrepancy between distributions of node representations [84, 68, 85, 174], graph representations [67], or labels [69].
One widely used distance metric $D$ in DRO for graphs is the Wasserstein distance [68, 69, 85]. Specifically, Zhang, et al. [68] propose a graph learning framework comprising two models based on two separate assumptions regarding data uncertainty. One model assumes that every distribution within the uncertainty set follows a Gaussian distribution, while the other operates without any priors. In another work [174], the uncertainty set is defined in a Laplacian form. Besides, based on the assumption that nodes belonging to the same class follow the same underlying feature distribution, the Wasserstein distance is used to define the uncertainty set that includes the underlying distribution of node embeddings for each class [69].
Furthermore, some work [84, 67] explore the connection between DRO (using the KL-divergence as the distance metric) and graph learning methods. Specifically, Wu et al. [67] reveal that contrastive learning implicitly performs DRO over the negative sampling distribution. Then they refine the “worst-case distribution” and assign new weights to negative samples (in contrastive learning). In another work [84], LightGCN [^3] LightGCN only includes the neighborhood aggregation component, and excludes the feature transformation and nonlinear activation components. [189] is proven to work as a graph smoothness regularizer within the DRO framework, enabling a regularization-based method that promotes model robustness and generalization.
V-B2 Adversasrial Training
Adversarial Training Overview. Adversarial Training (AT) first generates adversarial samples, which aim to manipulate the original predictions and degrade model performance most with imperceptible perturbations. The models will then be trained with these adversarial samples. Formally, AT is formulated as the following min-max optimization problem:
$\min_{\theta}\max_{\|\epsilon\|\leq\rho}\mathcal{L}_{0}(\theta,x+\epsilon,y).$ (33)
Adversarial Training for Graphs. Considering both the attributional and topological features of graphs, AT for graphs considers the following min-max optimization problem:
$\min_{\boldsymbol{\theta}}\max_{\epsilon_{\theta},\epsilon_{x},\epsilon_{a}}\mathcal{L}_{0}(\boldsymbol{\theta}+\epsilon_{\theta},X+\epsilon_{x},A+\epsilon_{a},Y),$ (34)
where $\epsilon_{\theta}$, $\epsilon_{x}$, and $\epsilon_{a}$ represent the perturbations applied to model parameters $\boldsymbol{\theta}$, node feature matrix $X$, and adjacency matrix $A$, respectively. These perturbations are intended to be small enough and imperceptible, constrained by a budget such that $D(X+\epsilon_{x},X)\leq\rho_{x}$, with a slightly abuse of $D$. Similar constraints are applied to $\epsilon_{\theta}$ and $\epsilon_{a}$. Consequently, the uncertainties from model parameters, graph topologies, and node features are incorporated during the generation of adversarial samples. We summarize the existing studies from these three aspects.
A branch of AT studies on graphs focuses on manipulating model parameters $\boldsymbol{\theta}$ to improve generalization [86, 87], where the perturbations are evaluated by $L_{p}$ norm.
Some studies [175, 71, 176, 70, 88] focus on manipulating graph topologies, such as adding or deleting edges, corresponding to $\epsilon_{a}$ in Eq. (34). For the distance metric $D$, many works [175, 71, 176, 70] count the number of malicious edge modifications by the $L_{0}$ norm. In other work [88], discrete edge modifications are relaxed to continuous changes, evaluated using the $L_{1}$ norm.
Besides adding or deleting nodes from the graph, perturbing node features is a commonly explored strategy [72, 89, 90, 73], referring to $\epsilon_{x}$ in Eq. (34). Most relevant works [72, 89, 90] adopt the $L_{p}$ norm (e.g., $L_{1}$, $L_{2}$, or $L_{\infty}$) to constrain perturbations, and the Wasserstein distance has also been sparsely used [73].
Unlike DRO methods, which use distributional distance to set a boundary between certain and uncertain data, most AT studies rely on $L_{p}$ norm as distance metrics to capture potential variations in test data. While the $L_{p}$ norm could measure the magnitude of perturbations, it is not designed to quantify uncertainty. Instead, the $L_{p}$ norm defines a bounded region around an input, indicating the extent of perturbation, which implies the epistemic uncertainty of the model has about the data. Moreover, as previously mentioned studies [68, 69, 174], DRO frameworks leverage uncertainty sets to quantify epistemic uncertainty by measuring how much the model’s performance varies across these uncertainty sets. These approaches make it possible to model the knowledge gap the model has about unseen or shifted data distributions. However, most AT works focus on making models robust to data shift via adversarial perturbation, lacking direct uncertainty quantification studies by AT.
V-B3 Self-Supervised Learning
Self-Supervised Learning Overview. Self-Supervised Learning (SSL) allows models to learn informative signals from input samples and their variants, reducing the request for labeled samples, and thus alleviating the need for expensive annotations [177, 178]. SSL methods typically design pretext tasks that enable models to learn generalized representations from unlabeled data, providing a good initialization for downstream tasks. Various pretext tasks are proposed to extract meaningful features from data.
Taking SSL on computer vision as an example, the pretext tasks include image inpainting, where models learn to predict the missing parts masked by design [179]; image colonizing, where models learn to colorize grayscale images [180]; jigsaw puzzle solving, where models learn image representations by arranging scrambled image patches [181]; and self-supervised contrastive learning, which leverages image transformations to enhance feature learning [182, 183]. Besides, SSL methods have also seen remarkable success in the natural language processing domain [184, 185, 186].
Self-Supervised Learning for Graphs. SSL on graphs involves deliberately perturbing or modifying graphs to construct new controllable datasets while preserving the intrinsic information of the original graphs. These created variations of the graph instances allow SSL to simulate real-world shifts in data distributions. Thus, the model can better handle epistemic uncertainty by training on these augmented graphs. We categorize the SSL methods based on designing structural and attributional features.
Some studies focus on altering the structural features of graphs. For instance, You et al. [91] propose augmentations where a portion of nodes is randomly removed, and a set of edges can be either deleted or added. These modifications generate new graph instances while maintaining the majority of global structure. Additionally, they propose subgraph extraction using random walk, which is supposed to preserve the local structure of the original graph. The subgraph extraction augmentation method is also employed in [74] to capture local structure information effectively.
Other studies propose to manually design modification for the attributional features of graphs. Specifically, inspired by image inpainting, graph completion is proposed in [92], where the attributional features of nodes are masked, and the model learns to recover these missing features. Similarly, Hu et al. [75] propose randomly masking attributional features on nodes and/or edges to create variations in the graph without changing its overall structure. Furthermore, You et al. [92] propose node clustering and graph partitioning methods, where nodes are clustered based on feature similarity or topological connections, and cluster indices are used as pseudo-labels for SSL tasks. Some works, like [93], augment graph data at both the structural and attribute levels, providing a more comprehensive approach to generating new graph views.
The objectives of SSL on graphs can vary significantly. For example, Yehudai et al. [187] propose a novel SSL task centered on predicting the d-pattern tree descriptor, which captures specific topological patterns within the graph. Meanwhile, Liu et al. [188] address the distribution shift issue caused by high-confidence unlabeled nodes. They propose a collaborative framework based on information gain to reweight samples, thereby mitigating the effects of this shift. This work highlights how SSL can adapt to changes in data distribution and reduce uncertainty for OOD samples.
Like AT, SSL studies often incorporate useful augmentations for robustness. While these augmentations help simulate different data variations and implicitly introduce uncertainty, they do not explicitly measure or quantify the epistemic uncertainty caused by these transformations. Nonetheless, SSL augmentations are often more diverse and challenging due to significant perturbations over inputs. It allows SSL to better generalize models across a wider range of possible variations in graph data.
V-C Calibration
Calibration Overview.
Definition of Calibration. We begin with a classifier $f$ trained by a feature set and a corresponding label set whose label space is $\mathcal{Y}=\{1,...,K\}$. With inputs $x$ and the true label $y$, which can be represented as a one-hot vector $\textbf{y}=(y^{1},...,y^{K})$, $f$ applies softmax operation to output a probability vector, $\textbf{v}=(v^{1},...,v^{K}),\sum_{j=1}^{K}v^{j}=1$. The predicted label and confidence of $x$, denoted by $\hat{y}$ and $\hat{p}$, respectively, are
$\hat{y}=\arg\max_{j\in\mathcal{Y}}v^{j},\text{ }\hat{p}=\max_{j\in\mathcal{Y}}v^{j}.$ (35)
A model is considered well-calibrated if the confidence matches the probability of correct prediction [9]: $\text{Pr}(\hat{y}={y}|\hat{p}=p)=p,\forall p\in[0,1]$. A stronger definition of being well-calibrated requires that each class (not only the predicted label) [95, 190] are calibrated. More formally, $\text{Pr}(y=j|v^{j}=p)=p,\text{ }\forall j\in\mathcal{Y},\text{ }\forall p\in[0,1]$.
However, the calibration performance of modern neural networks has been proven poor along a series of factors, including network width, depth, batch normalization, and weight decay [9].
Calibration Methods. Calibration can be conducted on top of a trained accuracy-driven model to adjust prediction confidences via a calibration set $\{(x_{i},y_{i})\}_{i=1}^{n}$. In this case, calibration methods are called post-processing or post-hoc [94, 95, 191, 96, 192]. A popular non-parametric calibration approach within this group is Histogram Binning [94]. The idea is to group predictions into bins according to their confidence values. For instance, in a binary classification task with label space $\mathcal{Y}=\{0,1\}$, the probability vector of $x_{i}$ is $\textbf{v}_{i}=(v_{i}^{0},v_{i}^{1})$. Each bin $B_{m}$ is a set including samples whose $v_{i}^{1}$ is in the interval $I_{m}=(\frac{m-1}{M},\frac{m}{M}]$ based on a user-defined $M$. Therefore, we can use a calibration set to minimize
where $\theta_{m}$ is the estimated number of positive-class samples in the $B_{m}$ set. For a test sample $x_{n+1}$ whose $v_{n+1}^{1}$ belong to $I_{m}$, $v_{n+1}^{1}$ should be replaced by $\theta_{m}$ as a calibrated confidence. Based on Histogram Binning, isotonic regression further optimizes the number and size of $I_{m}$ [95]. It is applied to deep learning models for regression tasks with promising performance [191]. Histogram Binning and isotonic regression can be applied to multi-class classification tasks using the one-vs-rest strategy. Nevertheless, both methods involve the creation of bins based on predicted confidence levels rather than directly on the input features, so the prediction’s confidence adjustment is determined by the bin to which it belongs, rather than by the characteristics of the input features themselves. As a result, neither method is adaptive to the input. In contrast, Platt scaling is a parametric input-adaptive calibration method [96, 193]. Traditionally, $\textbf{v}_{i}$ is transformed from a logit vector $\textbf{z}_{i}$ via softmax function $\sigma(\cdot)$. Platt scaling generalizes the process by $\textbf{v}_{i}=\sigma(\textbf{W}\cdot\textbf{z}_{i}+\textbf{b})$, which can be optimized by minimizing negative log-likelihood (NLL) on a calibration set. When W is a scalar and $\textbf{b}=0$, it is also referred to as temperature scaling. Accuracy versus uncertainty calibration (AvUC) loss [192] extends the temperature scaling method to encourage high confidence in accurate predictions and low confidence for likely incorrect predictions.
Some calibration techniques can be integrated directly into the training process, without requiring an additional calibration dataset [98, 97, 194]. With a training set $\{(x_{i},y_{i})\}_{i=1}^{n}$, a model is usually trained for a classification task NLL. However, NLL is minimized when the model assigns high probability to the correct class for each instance. In other words, the loss encourages the model to be confident in its predictions. While this is desirable when predictions are correct, it can lead to overconfident outputs even on uncertain or incorrect inputs. As a result, minimizing NLL too aggressively can lead to miscalibrated models that are overconfident and poorly represent uncertainty [194]. To mitigate the issue, label smoothing tunes a parameter $\alpha\in[0,1]$, so the predicted one-hot vector of $x_{i}$, denoted by $\textbf{y}_{i}=(y_{i}^{1},...,y_{i}^{K})$, becomes smoother by
$\mathbf{y}_{i}\rightarrow\mathbf{y^{\prime}}_{i}=\left((1-\alpha)y_{i}^{1}+\frac{\alpha}{K},...,(1-\alpha)y_{i}^{K}+\frac{\alpha}{K}\right),$ (37)
so it maps $\mathbf{y}_{i}$ to $\mathbf{y^{\prime}}$ by not assigning a full probability to the ground truth label [97]. Similarly, focal loss (FL) [98] is applied to calibration by
$\text{FL}(x_{i})=-(1-\hat{p}_{i})^{\gamma}\log(\hat{p}_{i}),\gamma\geq 0,$ (38)
where $\hat{p}_{i}$ is the prediction confidence of input $x_{i}$ [194]. Focal loss only considers confidence in the true label and the loss value will decrease faster as we increase $\gamma$, so it can prevent being over-confident. If $\gamma=0$, FL recovers CE.
Calibration for Graphs. Calibration on classification GNNs started to draw people’s attention recently as well [195, 99, 100, 101, 102, 130, 129, 103, 104, 37]. The works of calibrating GNNs can be categorized into node-level-classification, link prediction, and graph-level classification.
Node-level Classification. Traditional calibration techniques are usually applied to models handling independent and identically distributed (i.i.d.) data. In contrast, graph data presents a unique challenge due to its dependent nature. In graph data, the characteristics or status of one node may be influenced by the attributes or conditions of neighboring nodes, highlighting an inherent dependency where a node’s properties are not solely its own but also a reflection of its position within the graph [195]. This interdependency complicates the calibration process for GNNs. Furthermore, issues with class imbalance can lead to collapsed predictions, necessitating the adjustment of class weights during the training phase.
The calibration performance of GNNs on node-level classification tasks is related to GNNs’ architectures, neighbors’ labels, and the number of training epochs. Studies have shown that GNNs tend to produce underconfident predictions [99, 100]. Besides, it is proved that GNNs with wider layers tend to be better calibrated, and GNNs change from being under-confident to overconfident as the number of network layers increases [101, 102]. Calibration error in the classification of a node can be influenced by the node’s neighbor as well. Experimental results show that well-calibrated confidence distributions in graphs are often homophilic [99]—that is, neighboring nodes tend to have similar ground-truth confidences. Furthermore, it has been shown that calibration error decreases as the number of same-label neighbors increases [100], suggesting that strong local label agreement improves calibration. However, Liu et al. [101] report a contrasting observation: well-calibrated predictions are more likely when the same-class neighbor ratio is moderate (around 0.4 to 0.5), and models tend to become overconfident when the same-class neighbor ratio is too high. This apparent conflict between favoring high versus moderate same-label neighbor ratios may stem from differences in graph structures, model architectures, or training regimes (e.g., number of epochs). Besides, Liu et al. [101] show that GNNs tend to be under-confident with early stopping and will be overconfident with too many training epochs.
Most uncertainty quantification methods designed for multi-layer perceptrons can not recalibrate GNNs well. Teixeira et al.[195] and Vos et al.[196] show that methods like histogram binning, isotonic regression, Platt scaling, Accuracy-versus-Uncertainty (AvU)[197], and Graph Calibration Loss (GCL)[102]—which will be introduced shortly—generally outperform approaches such as Monte Carlo dropout [171] and model ensembles [198]. While Monte Carlo dropout and ensembles can improve calibration by capturing model uncertainty through stochastic regularization and diversity across multiple models, respectively, their effectiveness is often limited by computational cost and the lack of post-hoc adjustment. In contrast, the former methods directly refine the confidence estimates, leading to better-calibrated predictions. Moreover, models optimized on multi-class node classification tasks will be better calibrated and more discriminative than those trained on binary classification tasks [196]. Based on these insights, some works developed calibration approaches customized to GNNs.
The calibration methods for GNNs can be post hoc [99, 100, 101]. CaGCN is a topology-aware post-hoc calibration GCN model assuming that neighboring nodes share similar ground truth confidence levels with the center node [99]. With an undirected attributed GCN optimized for prediction accuracy, CaGCN works as another GCN to calibrate the outputs of the accuracy-driven model. Hence, CaGCN can be regarded as a Platt scaling method adapted to graph models. To preserve the accuracy of the original classification GCN, the optimization objective function of CaGCN is a combination of negative log-likelihood loss and a customized calibration loss, which aims to calibrate under-confident correct predictions and over-confident incorrect predictions. Graph Attention Temperature Scaling (GATS) [100] is also a post-hoc method, and embeds the attention mechanism for node-wise calibration. For each node, a scaling process will consider global calibration bias, relative confidence compared with neighbors, and aggregated neighborhood influence with corresponding attention coefficients. Ratio-binned scaling (RBS) follows the logic of histogram binning to calibrate outputs of GNNs [101]. RBS first estimates the same-class-neighbor ratio of each node, then groups the nodes into bins according to their ratio values, and finally learns a temperature for each bin on a validation set.
Conducting calibration during model training is also attractive since it does not require additional data and computation costs. Graph Calibration Loss (GCL) enables an end-to-end calibration method for GNNs [102]. GCL minimizes the discrepancy between the ground-truth and predicted confidence distributions. Specifically, GCL is the KL divergence between the two distributions regularized by the product of their entropies. HyperU-GCN is proposed to calibrate the hyperparameter uncertainty of automated GNNs using bi-level optimization [199]. The upper-level problem explains uncountable uncertainties by developing probabilistic hypernetworks through a variational Bayesian method. The lower-level problem learns how the hyperparameter distribution is propagated to GCN weights.
To maintain well-calibrated predictions in the presence of adversarial attacks and discriminatory biases, Multi-view Confidence-calibrated NIFTY (MCCNIFTY) was introduced [129]. Previous methods, such as FairGNN [200], RobustGNN [201], and NIFTY [202], primarily focused on preserving accuracy under such conditions, but often produced overconfident incorrect predictions or were vulnerable to global graph noise. MCCNIFTY combines intra-view calibration, inter-view fusion, and uncertainty-aware message passing to generate well-calibrated predictions under the impact of adversarial perturbations and biases.
Link Prediction. Link prediction is an important downstream task of GNNs [37]. Recent research finds that GNNs tend to be overconfident in negative link predictions but underconfident in positive ones [130].
IN-N-OUT [130] calibrates link prediction by learning a temperature parameter through edge perturbations. In this approach, the embedding of an edge in a GNN depends on the predicted labels of its two connected nodes. During training, two distinct embeddings are computed for each edge: one with the edge present in the graph (included in message passing) and one with the edge removed (excluded from message passing). The discrepancy between these two embeddings is utilized to determine the appropriate value for temperature scaling. For example, a high discrepancy between the two embeddings suggests that the confidence associated with the existence of the edge should be reduced.
In contrast, E-$\Delta$UQ [103] calibrates link prediction by perturbing node features. It is based on an anchoring technique, called $\Delta$-UQ [203]. Specifically, $\Delta$-UQ randomly chooses an anchor $c$, which is a constant bias with the same size of input feature, for stochastic centering during training, and creates a relative representation $[x-c,c]$ of each input $x$. Therefore, a model is trained on $([x-c,c],y)$ instead of $(x,y)$. The anchor value $c$ is different in each training iteration. During inference, $\Delta$-UQ generates multiple predictions for an input $x$ using different random perturbations $c$. If the model is confident about a particular label, the corresponding softmax scores remain stable across these random anchoring. In contrast, uncertain predictions exhibit higher variability. The final calibrated confidence distribution over labels is obtained by subtracting the standard deviation from the mean of the predicted softmax distributions across perturbations. The calibrated output of input $x$ is the label with the highest value in this adjusted confidence distribution. E-$\Delta$UQ extends the method to calibrate link prediction through two stages: encoding node information and decoding to predict links between node pairs. Random anchoring is applied during the encoding and/or decoding stages to enable uncertainty-aware training. It has three variants. The first variant implements anchoring between encoding and decoding, thus it is based on a deterministic encoder but the decoder is stochastic. The second variant implements one more anchoring after specific layers of encoding. Based on the second one, the last variant pre-trains the encoder with an auxiliary attribute masking task on node features. Although both IN-N-OUT and E-$\Delta$UQ rely on perturbation, E-$\Delta$UQ offers greater flexibility since its anchoring is not limited to binary perturbations as in IN-N-OUT.
Graph-level Classification. Under the conditions of covariate and concept shifts, the anchoring technique is also adopted in G-$\Delta$UQ [104] with three variants for the calibration of GNNs on graph-level classification using Superpixel-MNIST datasets [204, 205, 206, 207]. The first variant, node feature anchoring, applies stochastic centering directly to the input node features. The second, intermediate anchoring, introduces stochastic centering after specified hidden layers for each node. The final variant, readout anchoring, applies anchoring to the overall graph representation before the graph-level readout function, rather than to individual nodes.
V-D Conformal Prediction
Conformal Prediction Overview. Given a test input, conformal prediction (CP) generates a prediction set (rather than a single prediction) based on a statistical threshold derived from uncalibrated confidences of previous predictions. The prediction set guarantees the inclusion of the true label with a specified probability [105] or an upper bound of a user-defined prediction loss [208]. We concentrate on the use of CP in classification tasks, as GNNs are commonly employed in this context.